OpenStudy (anonymous):

a pyramidal tent has a square base of side 4m. the height of each triangular face is 6 m. what is the area of the material which covers all the sides of the tent?

OpenStudy (aripotta):

including the base?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

uhmm. wait.

OpenStudy (aripotta):

we'll just do without the base

OpenStudy (anonymous):



OpenStudy (anonymous):

square based pyramid..

OpenStudy (aripotta):

LA pyramid = pl/2 where p means perimeter of base and l means slant height. perimeter of base = 4 + 4 +4 +4 = 16 16(6)/2 = 96/2 = 48m^2